OpenStudy (anonymous):

Bob has 10 more marbles than Larry. Together they have a total of 29 marbles. How many marbles does Larry have? Let L = the number of marbles Larry has.

OpenStudy (anonymous):

The number of marbles Bob has is L + 10 SO total: L + 10 + L = 29 2L + 10 = 29 2L = 19

OpenStudy (anonymous):

But then I end up with L = 9.5... I think I did something wrongly/ there's a problem with the question.. Sorry, not able to help :/

OpenStudy (anonymous):

Easy: 29-10=19 L=29-10

OpenStudy (anonymous):

@MathCrazyFool's logic is correct...check the original problem and make sure the numbers are correct. @saraneedshelplease is incorrect...if Larry has 19 marbles then Bob has 19 + 10 = 29 and together they have 19 + 29 = 48.

OpenStudy (anonymous):

I just said "29-10" which would be 19. Larry has 19 marbles. If Larry had 19 marbles and Bob had 10 more than Larry: 19+10=29. @ChristianGeek

OpenStudy (anonymous):

@saraneedshelplease read the problem again...it says that TOGETHER they have a total of 29 marbles, not Bob has 29 marbles.